160 mm + B 480 N 60 mm 80 mm # 150 mm 60 mm 100 mm

For the frame shown find the force in link DE and forces at C. I know the answers but I need the work. The Force in DE is 561 N, Cx is 795 N left and Cy is 216 N up.
Answer #1

A My Ax Omrin 480N & Ey zo it Ay – 480 =0 > Ay = 480N EMA=0 9+ (RB) Cibomm) – (4801) (100mm) 20 C RB = 300 N Free body diagraEMczo u + E300N) (60) - (480N) Cloomm) – (Fco Susa) 250mm =0 6300N) (60) - (480) (100) - (Fepsió 28.07)(250) =0 FD=-561N Fo=5